OpenStudy (anonymous):

What are the coordinates of the center of the ellipse shown below? my answer is (7,-3)

OpenStudy (anonymous):

\[\frac{ (x-7)^2 }{ 4 } + \frac{ (y+3)^2 }{ 16 } = 1 \]

OpenStudy (anonymous):

critical situation lol soo i must be 100% sure hahah thats why i am double checking ..

OpenStudy (anonymous):

its correct :)

OpenStudy (anonymous):

Yay thank you